How many quarts of a 10% anti-freeze solution must be mixed with 16 quarts of a 40% anti-freeze solution to make a 30% solution?
Answer by stanbon(75887)

How many quarts of a 10% anti-freeze solution must be mixed with 16 quarts of a 40% anti-freeze solution to make a 30% solution?
------------------------------
Anti-freeze equation:
afreeze + afreeze = afreeze
0.10x + 0.40*16 = 0.30(x+16)
----------
Multiply thru by 100 to get:
10x + 40*16 = 30x + 30*16
20x = 10*16
x = 8 quarts (amt. of 10% solution needed for the mixtue)
=================================
Cheers,
Stan H.